OpenStudy (mackenzie2013):

Twenty-seven is __ % of 60.

OpenStudy (anonymous):

45%

OpenStudy (anonymous):

100/60*27

OpenStudy (anonymous):

27=(x/100)*60 27/60=x/100 x=(27/60)*100 =(9/20)*100=45

OpenStudy (mackenzie2013):

So you just take 60 and divide it by100 and then multiply by 27?

OpenStudy (anonymous):

no divide 100 by 60 then multiply it by 27

OpenStudy (gorv):

divide 27 by 60 and multiply by 100

OpenStudy (uri):